a rectangle is three times as long as it is wide. its perimeter is 36. find each side
Answer by robertb(5830)

2(w + 3w) = 36,
2(4w) = 36,
8w = 36,
w = 36/8 = 9/2, the width of the rectangle.
l = 3w = 27/2, the length of the rectangle.